Consider the following demand function: QD = 25/P2. Show that the point elasticity of demand for that function is always equal to −2.


1
Expert's answer
2022-02-17T04:37:51-0500

"E_d= \frac{\Delta Q}{\Delta P}\times \frac{P}{Q}"

"\frac{\Delta Q}{\Delta P}= Slope of the demand function"

"Q_D= \frac{25}{P^2}"

"Q_D= 25P^{-2}"

"\frac{\Delta Q}{\Delta P}= -50P^{-3}"